OpenStudy (anonymous):

Karen opens a savings account with $1500. She deposits $100 every month into the account that has a 0.85% interest rate, compounded annually. If she doesn’t withdraw any money, what will the account balance be in 10 years? $14,102.05 $14,208.04 $12,575.55 $12,469.56

OpenStudy (anonymous):

Have you tried solving this yourself? It's quite a big problem to just copy and paste on here. I'd like to know what you did, where you think you went wrong, etc.

OpenStudy (anonymous):

I used the future value formula and the compound interest formula

OpenStudy (anonymous):

im not getting any of the answers

OpenStudy (anonymous):

what would be the present value?

OpenStudy (anonymous):

so what would be the formula?

OpenStudy (phi):

this is the same problem as yesterday

OpenStudy (phi):

the 2nd problem in this thread http://openstudy.com/users/chelseachels#/updates/51487480e4b05e69bfaaef33

OpenStudy (anonymous):

the PMT(((1+I)^n)-1)?

OpenStudy (phi):

yes

OpenStudy (anonymous):

would the payment be $1500 or $100?

OpenStudy (phi):

however, because it is compounded annually but payments are made monthly it is tricky

OpenStudy (anonymous):

so then what do i do?

OpenStudy (phi):

good question. I have not seen this problem before, and would have to figure it out. However, I would expect that your book tells you how to do it. Can you find a similar problem in your book and post how they solve it ?

OpenStudy (anonymous):

im in flvs so there is no book and the examples dont show how to do this problem

OpenStudy (phi):

I am sure you have been taught how to do this... But I'll post something later on this problem.

OpenStudy (anonymous):

ok thanks

OpenStudy (mertsj):

In the meantime, you could grind it out one year at a time as I did and at the end of 10 years, I got the second choice.

OpenStudy (anonymous):

can you write on here what steps you took?

OpenStudy (mertsj):

There is no interest paid until the end of the first year at which time there will be 1500+1200 in the account. So multiply 2700 by 1.0085 to get the value of the account at the end of the first year.

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

So now there is 2722.95 in the account. Interest will not be paid again until the end of the second year and at that time there will be 3922.95 in the account. Multiply that by 1.0085 to get 3956.30 at the end of the second year.

OpenStudy (anonymous):

how did you get 3956.30?

OpenStudy (mertsj):

Add 1200...Multiply by 1.0085 Until you have reached the end of the 10th year.

OpenStudy (mertsj):

at that time there will be 3922.95 in the account. Multiply that by 1.0085

OpenStudy (anonymous):

oh ok

OpenStudy (mertsj):

3922.95 times 1.0085

OpenStudy (anonymous):

3956.30 i understand that

OpenStudy (anonymous):

and then you keep going until you get to the tenth year?

OpenStudy (mertsj):

Add 1200...Multiply by 1.0085 Until you have reached the end of the 10th year.

OpenStudy (anonymous):

thanks so much

OpenStudy (mertsj):

yw

OpenStudy (phi):

Here is what I came up with: opens a savings account with $1500. after 10 years, at interest rate 0.0085/year compounded yearly P1= 1500* (1+0.0085)^10 = 1632.49 for the 2nd part, 100 per month, for 10 years The trick (for me) is that you do not get interest until the end of the year, so after 12 months, she has put in 1200. now we find the interest. This is like the problem of making a payment once a year of 1200, That means we can use the formula \[ PMT\frac{ (1+i)^n -1}{i} \] PMT is 1200 I is 0.0085 n is 10 (years) we get : 1200* ( (1.0085)^10 -1)/0.0085 = 12469.56 (notice that is one of the choices) but we have to add in the dollars from the initial 1500: 1632.49 total is 1632.49 + 12469.56 = 14102.05